### v1

# Incorporating Damped Harmonic Oscillator in DSGE Models

Incorporating Damped Harmonic Oscillator in DSGE Models
Wei Chun Hsu
(February 10, 2025)
1 Introduction
Economic fluctuations, commonly referred to as business cycles, have long been a central concern for macroeconomists. These fluctuations arise due to various internal and external factors, including technological changes, monetary policy shifts, and international economic events. As economies face these shocks, understanding how they adjust and recover becomes critical for policymakers, businesses, and households.
Traditionally, Dynamic Stochastic General Equilibrium (DSGE) models have provided a robust framework for analyzing these fluctuations. DSGE models derive macroeconomic outcomes from the optimizing behavior of individual agents, such as consumers and firms, who respond to shocks under rational expectations. However, DSGE models are often criticized for their inability to fully capture delayed adjustments seen in real-world economies, where frictions such as price and wage stickiness slow down the return to equilibrium.
To address these shortcomings, we turn to physics, specifically the concept of the damped harmonic oscillator. This model describes systems where oscillations gradually decay due to friction or resistance, eventually leading to a stable equilibrium. The damped harmonic oscillator’s ability to model systems that experience delayed or slow responses makes it a valuable tool for improving DSGE models. By incorporating a damping coefficient, we can model economies that exhibit varying recovery speeds, depending on the extent of market frictions and policy delays.
This paper aims to integrate the damped harmonic oscillator into DSGE models to create a more realistic framework for analyzing economic fluctuations. We explore how different levels of damping—under-damped, critically damped, and over-damped—affect the speed and stability of economic recovery. Through numerical simulations and historical case studies, we demonstrate the model’s applicability in explaining real-world economic events, such as the Great Depression and the 2008 Global Financial Crisis.
2 Literature Review
2.1 Evolution and Application of DSGE Models
The development of DSGE models has its roots in Real Business Cycle (RBC) theory, introduced by Kydland and Prescott (1982). RBC models focus on technological shocks as the primary drivers of economic fluctuations, treating recessions and booms as natural responses to changes in productivity. While RBC models provide a solid microeconomic foundation, they fail to account for market imperfections, such as price stickiness or wage rigidity, which are critical for understanding short-term economic dynamics.
The New Keynesian DSGE models addressed many of these limitations by incorporating nominal rigidities, such as sticky prices and wages. These models allow for the analysis of monetary policy’s short-term impact on output and employment, particularly in the presence of demand shocks. For instance, Woodford (2003) highlights the role of central banks in stabilizing inflation and output using interest rate rules within DSGE models.
However, even New Keynesian DSGE models tend to assume a relatively quick adjustment to equilibrium following shocks. This assumption underestimates the persistence of economic slumps, especially when significant market frictions or structural issues are present. These models do not adequately capture the delayed responses seen in real-world recoveries, where economies can experience prolonged periods of stagnation or instability before returning to normal levels.
2.2 Damped Harmonic Oscillator: Basic Theory and Mathematical Structure
The damped harmonic oscillator is a classical model in physics that describes systems where an external force causes oscillations, which are gradually reduced by internal resistance or damping. The system is characterized by the following differential equation:
$m\ddot{x}+c\dot{x}+kx=F(t)$
Where:
  • $m$ is the mass of the object,
  • $c$ is the damping coefficient (friction),
  • $k$ is the spring constant (restoring force),
  • $F(t)$ is the external force acting on the system.
The behavior of the system depends on the magnitude of the damping coefficient $c$, which determines how quickly the system’s oscillations decay:
  • Under-damped ($c^{2}<4mk$): The system oscillates with gradually decreasing amplitude.
  • Critically-damped ($c^{2}=4mk$): The system returns to equilibrium without oscillating and in the shortest possible time.
  • Over-damped ($c^{2}>4mk$): The system returns to equilibrium slowly without oscillations.
This dynamic behavior is similar to how economies adjust after a shock. In an economic context, the damping coefficient $\gamma$ can represent various frictions, such as price stickiness, wage rigidity, or policy delays, that slow down recovery. Introducing this concept into DSGE models allows for a more accurate representation of how economies recover from shocks.
2.3 Cross-disciplinary Research between Economics and Physics
The field of econophysics has seen an increasing amount of research in recent decades. Econophysics applies models and methods from statistical physics to economic phenomena, particularly in financial markets. For example, Bouchaud and Potters (2003) used statistical physics to analyze market volatility and asset pricing.
However, while econophysics has primarily focused on financial markets, there has been less research on applying physical models like the damped harmonic oscillator to macroeconomic fluctuations. This paper builds on previous work by integrating the mathematical framework of the damped harmonic oscillator into DSGE models, thereby creating a novel approach to understanding economic recovery dynamics.
3 Theoretical Foundation
3.1 Basic Structure of DSGE Models
DSGE models are grounded in the optimizing behavior of individual economic agents. Consumers maximize their utility over time, typically represented by an intertemporal utility function:
$\max\mathbb{E}_{0}\sum_{t=0}^{\infty}\beta^{t}U(C_{t},L_{t})$
Where:
  • $C_{t}$ is consumption at time $t$,
  • $L_{t}$ is leisure,
  • $\beta$ is the discount factor, representing the trade-off between present and future consumption.
Households make decisions subject to a budget constraint:
$C_{t}+B_{t+1}=(1+r_{t})B_{t}+W_{t}N_{t}$
Where:
  • $B_{t}$ is bond holdings,
  • $r_{t}$ is the interest rate,
  • $W_{t}$ is the wage rate,
  • $N_{t}$ is labor supply.
On the production side, firms maximize profits by choosing labor $N_{t}$ and capital $K_{t}$ inputs to produce output:
$\max\Pi_{t}=P_{t}Y_{t}-W_{t}N_{t}-R_{t}K_{t}$
Where $Y_{t}$ is the firm’s output, $P_{t}$ is the price level, $W_{t}$ is wages, and $R_{t}$ is the rental rate of capital. Firms face a production function:
$Y_{t}=A_{t}F(K_{t},N_{t})$
Where $A_{t}$ represents total factor productivity (TFP).
Market-clearing conditions ensure that the sum of supply and demand in both goods and labor markets equals zero. These individual optimization problems and market-clearing conditions are combined to derive a general equilibrium.
3.2 Dynamics of the Damped Harmonic Oscillator
The damped harmonic oscillator is a useful framework for describing the delayed response of an economic system to external shocks. In this model, the output gap $Y_{t}$ represents the deviation of actual output from potential output. The dynamic equation of the system is given by:
$\ddot{Y}_{t}+\gamma\dot{Y}_{t}+\alpha Y_{t}=\varepsilon_{t}$
Where:
  • $\gamma$ is the damping coefficient, representing market frictions or delays in policy implementation,
  • $\alpha$ is the system’s natural adjustment frequency, which dictates how fast the economy would return to equilibrium in the absence of frictions,
  • $\varepsilon_{t}$ is a stochastic shock representing external disturbances, such as financial crises or policy shocks.
The behavior of the system depends on the value of the damping coefficient $\gamma$:
  • Under-damped: If $\gamma$ is small, the system will exhibit oscillatory behavior, where the economy experiences fluctuations around equilibrium before eventually stabilizing.
  • Critically-damped: If $\gamma$ reaches a critical value, the system will return to equilibrium in the shortest possible time without oscillating.
  • Over-damped: If $\gamma$ is large, the economy will return to equilibrium slowly, without oscillations, reflecting a more gradual recovery process.
4 Mathematical Model Construction
4.1 Mathematical Representation of DSGE Models
In the standard DSGE framework, the economy’s dynamic behavior is driven by the intertemporal choices of households and firms. The Euler equation represents the optimal intertemporal choice of consumption:
$U^{\prime}(C_{t})=\beta(1+r_{t+1})\mathbb{E}_{t}U^{\prime}(C_{t+1})$
This equation shows that the marginal utility of consumption today must equal the expected discounted marginal utility of future consumption, adjusted for the interest rate. Firms’ investment decisions are driven by a similar intertemporal optimization problem, and the Tobin’s Q theory of investment links current capital accumulation to future profits.
5 Numerical Simulations and Estimating Damping Coefficient
5.1 Numerical Methods for Solving the Model
In this section, we simulate the dynamic behavior of the output gap $Y_{t}$ under various damping scenarios. The model used is based on the following second-order differential equation:
$\ddot{Y}_{t}+\gamma\dot{Y}_{t}+\alpha Y_{t}=\varepsilon_{t}$
Here, $\gamma$ is the damping coefficient representing market friction or policy delays, $\alpha$ is the system’s natural frequency, and $\varepsilon_{t}$ represents stochastic shocks. To solve this equation numerically, we employ the **Euler method**. This method approximates the solution of differential equations by iterating over small time steps and updating the values of $Y_{t}$ and $\dot{Y}_{t}$.
The following R code demonstrates the implementation of the Euler method:
# R code for simulating output gap dynamics under different damping conditions
library(ggplot2)
alpha <- 1.0  # Natural frequency of adjustment
time <- seq(0, 20, by=0.1)  # Time steps
dt <- time[2] - time[1]  # Time step size
epsilon <- 0  # External shock (set to zero for simplicity)
# Euler method for solving the ODE
solve_ode <- function(gamma, Y0, Ydot0, alpha, time) {
  Y <- numeric(length(time))
  Ydot <- numeric(length(time))
  # Initial conditions
  Y[1] <- Y0
  Ydot[1] <- Ydot0
  # Iterate over time steps using Euler method
  for (i in 2:length(time)) {
    Yddot <- -gamma * Ydot[i-1] - alpha * Y[i-1] + epsilon  # Acceleration term
    Ydot[i] <- Ydot[i-1] + Yddot * dt  # Update velocity
    Y[i] <- Y[i-1] + Ydot[i-1] * dt  # Update position
  return(Y)
# Initial conditions
Y0 <- 1  # Initial output gap
Ydot0 <- 0  # Initial velocity
# Damping coefficients for different scenarios
gamma_under <- 0.5  # Under-damped
gamma_critical <- 2.0  # Critically-damped
gamma_over <- 4.0  # Over-damped
# Simulate the output gap under different damping conditions
Y_under <- solve_ode(gamma_under, Y0, Ydot0, alpha, time)
Y_critical <- solve_ode(gamma_critical, Y0, Ydot0, alpha, time)
Y_over <- solve_ode(gamma_over, Y0, Ydot0, alpha, time)
# Combine data for plotting
data <- data.frame(
  Time = rep(time, 3),
  OutputGap = c(Y_under, Y_critical, Y_over),
  Damping = factor(rep(c("Under-damped", "Critically-damped", "Over-damped"), each=length(time)))
# Plot the results
ggplot(data, aes(x=Time, y=OutputGap, color=Damping)) +
  geom_line(size=1) +
  labs(title="Output Gap Dynamics under Different Damping Conditions",
       x="Time", y="Output Gap (Y)") +
  theme_minimal() +
  scale_color_manual(values=c("blue", "green", "red"))
This simulation helps us visualize how the output gap behaves under various damping scenarios.
5.2 Results and Interpretation
- **Under-damped** ($\gamma^{2}<4\alpha$): The system exhibits oscillations, indicating that the economy initially overshoots before stabilizing. This behavior is typically seen when policy interventions are too aggressive, leading to temporary volatility. - **Critically-damped** ($\gamma^{2}=4\alpha$): The system returns to equilibrium as quickly as possible without oscillating. This scenario is ideal, as it suggests a well-calibrated policy response. - **Over-damped** ($\gamma^{2}>4\alpha$): The system slowly converges to equilibrium without overshooting. This is often a result of cautious policies or significant market frictions, leading to prolonged recoveries.
5.3 Estimating the Damping Coefficient
To estimate the damping coefficient $\gamma$, we can apply statistical techniques such as Maximum Likelihood Estimation (MLE) or time-series models like Generalized Autoregressive Conditional Heteroskedasticity (GARCH). Below is an example of how to estimate $\gamma$ using MLE in R:
# Example observed data for output gap
Y_obs <- c(1, 0.8, 0.6, 0.5, 0.3, 0.2, 0.15, 0.1, 0.05)
# Negative log-likelihood function for MLE
nll <- function(gamma) {
  alpha <- 1  # Fixed natural frequency
  Y_pred <- numeric(length(Y_obs))
  Y_pred[1] <- 1  # Initial condition
  for (i in 2:length(Y_obs)) {
    Y_pred[i] <- Y_pred[i-1] * exp(-gamma * i)
  -sum(dnorm(Y_obs, mean = Y_pred, sd = 0.1, log = TRUE))
# Estimate damping coefficient gamma using MLE
mle_result <- mle(nll, start = list(gamma = 0.5))
summary(mle_result)
This approach estimates $\gamma$ by minimizing the negative log-likelihood function, comparing predicted and observed output gap values.
6 Application of Findings: Integrating with Financial Data and Theoretical Models
To apply the theoretical framework of damping in DSGE models with the empirical insights from the referenced literature, such as **Camera (2007)** and **Matsui (2008)** on credit systems and international currency, we need to understand the role of liquidity, policy interventions, and market frictions. The damping coefficient $\gamma$ can be directly interpreted as capturing the liquidity buffers or constraints imposed on the system, influencing both output fluctuations and the response to shocks.
For instance, in **Chang and Li (2016)**, liquidity effects due to central bank interventions provide key insights into how money injections alter the liquidity structure, which can be modeled through variations in the damping coefficient. The empirical analysis shows that unexpected money injections can have non-neutral effects on output and interest rates, depending on the liquidity conditions. By applying these findings to our model, we can simulate scenarios where $\gamma$ varies based on different levels of liquidity, offering a more nuanced view of policy effectiveness.
7 Conclusion
Incorporating damping effects into DSGE models allows for a more realistic representation of economic dynamics in the face of shocks. By simulating different damping conditions, we gain insights into how policies and market frictions influence the speed and stability of economic recoveries. Estimating the damping coefficient $\gamma$ using statistical methods further enriches the model by enabling empirical validation against real-world data.
The integration of theories on liquidity, credit systems, and international currency competition highlights the importance of understanding how different economic policies affect both the domestic and global financial landscape. Future research could extend this framework by examining sectoral differences and refining the model to capture additional factors influencing the damping coefficient.